OpenStudy (anonymous):

- 4/5ivided by 4/15 what about this

OpenStudy (anonymous):

- 4/5ivided by 4/15 what about this

OpenStudy (anonymous):

to divide by a fraction, multiply by the reciprocal

OpenStudy (anonymous):

-4/5 divided by 4/15 = what is that

OpenStudy (anonymous):

\[-\frac{4}{5}\times \frac{15}{4}\] cancel first, multiply last

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

please use divided as "/" and times as this "x" answer: (((-4) / 5) / 4) / 15 = -0.0133333333

OpenStudy (anonymous):

\[\frac{4\times 15}{4\times 3}\] \[\frac{\cancel{4}^1\cancel{15}^3}{\cancel{4}^1\cancel{3}^1}\]

OpenStudy (anonymous):

-4 over 5 divided by -4 over 15 sorry

OpenStudy (anonymous):

you should get -3

OpenStudy (anonymous):

in that case you should get 3

OpenStudy (anonymous):

your confusing him..her..

OpenStudy (anonymous):

You should find for the LCD of that two fractions for you to easily answer the fractions .

OpenStudy (anonymous):

hershey, you can't give decimal approximations to an answer that involves fractions

OpenStudy (anonymous):

i think the answer is 1/3 ?

OpenStudy (anonymous):

(-7)(-2)(-3)

OpenStudy (anonymous):

-42 ?